# 12, Proskura st., Kharkov 61085,Ukraine

### v1

@@ Abstract
where $\sigma_{3}=\mbox{diag}\left(1,-1\right)$? The answer, which we present here without derivation, is ‘yes’. However, to do this one should start not with the bright solitons of the ALH, as in section 3, but with the dark ones [18, 19]. In some sense the signature of the matrix describing the applied constraints play the role of the sign in front of the nonlinear term in the nonlinear Schrödinger equation: it determines which kind solitons (bright or dark) exists in the system.
Finally we would like to add a short comment on the integrability of the system (1.3) which was not discussed in the paper. We cannot at present prove its integrability, for example, by developing the inverse scattering transform. However, we now know that (1) it is a reduction of equations of the integrable ALH and that (2) it possesses $N$-soliton solutions. These two facts are a strong indication that (1.3) is integrable. Nevertheless, we think that the work in this direction should be continued, and one of the first problems to solve is to find the conservation laws of the model, which may be a topic of the following study.
Appendix.
@@
$\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}}-\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}$ $=$ $\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})|b\rangle-\langle{a}|\mathsf{R}^{-1}\mathsf{F}|b\rangle$
$=$ $\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})|b\rangle-\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}\mathsf{J}_{\xi}\mathsf{F}|b\rangle$
@@
$\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}}-\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}=\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{X}\mathsf{F}|b\rangle$ (A.1)
@@
$(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{X}\mathsf{F}=\mathbb{E}_{\xi}{\mathsf{F}}-\mathsf{J}_{\xi}\mathsf{F}.$
@@
$\mathsf{X}$ $=$ $1-\mathsf{J}_{\xi}+\mathsf{B}\mathsf{A}-(\mathbb{E}_{\xi}{\mathsf{B}})(\mathbb{E}_{\xi}{\mathsf{A}})\mathsf{J}_{\xi}$
@@
$=$ $\xi\mathsf{R}^{-1}\mathsf{F}^{-1}+\xi\mathsf{R}^{-1}|b\rangle\langle b|\mathsf{L}^{-1}\mathsf{A}.$
Substituting this expression into (A.1) and using (3.8) together with the identity $\mathsf{A}\mathsf{F}=\mathsf{G}\mathsf{A}$, one arrives at
$\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}}-\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}$ $=$ $\xi\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{R}^{-1}|b\rangle+\xi\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{R}^{-1}|b\rangle\langle b|\mathsf{L}^{-1}\mathsf{A}\mathsf{F}|b\rangle$
@@
$\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}}-\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}$ $=$ $\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})(\mathbb{E}_{\xi}{\mathsf{B}})|a\rangle-\langle a|\mathsf{R}^{-1}\mathsf{F}\mathsf{B}|a\rangle$ (A.2)
$=$ $\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{Y}\mathsf{G}|a\rangle$
@@
$(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{Y}\mathsf{G}=(\mathbb{E}_{\xi}{\mathsf{F}})(\mathbb{E}_{\xi}{\mathsf{B}})-\mathsf{J}_{\xi}\mathsf{F}\mathsf{B}.$
@@
$\mathsf{Y}$ $=$ $\mathbb{E}_{\xi}{\mathsf{B}}-\mathsf{J}_{\xi}\mathsf{B}+(\mathbb{E}_{\xi}{\mathsf{B}})\mathsf{A}\mathsf{B}-(\mathbb{E}_{\xi}{\mathsf{B}})(\mathbb{E}_{\xi}{\mathsf{A}})\mathsf{J}_{\xi}\mathsf{B}$
@@
$=$ $-\xi\mathsf{R}^{-1}|b\rangle\langle b|\mathsf{L}^{-1},$
and substituting it in (A.2) one can obtain
$\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}}-\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}$ $=$ $-\xi\langle b|\mathsf{L}^{-1}\mathsf{G}|a\rangle\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{R}^{-1}|b\rangle$

### v2

@@ Abstract
where $\sigma_{3}=\mbox{diag}\left(1,-1\right)$? The answer, which we present here without derivation, is ‘yes’. However, to do this one should start not with the bright solitons of the ALH, as in section 3, but with the dark ones. In some sense the signature of the matrix describing the applied constraints play the role of the sign in front of the nonlinear term in the NLSE: it determines which kind solitons (bright or dark) exists in the system.
Finally we would like to add a short comment on the integrability of the system (1.3) which was not discussed in the paper. We cannot at present prove its integrability, for example, by developing the inverse scattering transform. However, we now know that equations (1.3) possess $N$-soliton solutions. That means that the model (1.1) passes the so-called $N$-soliton test for integrability, which, though not proved rigorously, has a long history of successful applications and was even used as a tool for finding new integrable models, as, for example, in the comprehensive study by Hietarinta [9, 10, 11, 12]. Also, we now know that equations (1.3) are a consequence of equations of the integrable ALH. These two facts are a strong indication that the model (1.1) is integrable. Nevertheless, we think that the work in this direction should be continued, and one of the first problems to solve is to find the conservation laws of the model, which may be a topic of the following study.
Acknowledgments.
We would like to thank the Referees for careful reading of the manuscript and their valuable comments and suggestions, which helped to improve this paper.
Appendix A. A proof of (2.26).
To prove the fact that functions $u$ and $v$ satisfy the NLSE we need some consequences of the formulae presented in section 2.
First, it follows from equations (2.13) and (2.14) together with the restriction (2.10), that functions $u$ and $v$ can be written as
$u=i\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}_{x}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}-i\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}_{x},\qquad v=i\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}_{x}-i\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}_{x}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}$ (A.1)
and that
$uv=\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}_{x}\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}_{x}+\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}_{x}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}_{x}.$ (A.2)
Differentiating (A.1) with respect to $t$ and expressing the $t$-derivatives from (2.15) and (2.16) leads to
$\begin{array}[]{lcl}u_{t}&=&-u_{xx}+\left(\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}_{x}\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}+\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}_{x}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}\right)u_{x}\\ &&+\left(\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}_{xx}+\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}_{xx}-\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}_{x}\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}_{x}-\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}_{x}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}_{x}\right)u.\end{array}$ (A.3)
Another consequence of (2.13) and (2.14) is that
$\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}_{x}\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}+\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}_{x}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}=0,\qquad\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}_{x}+\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}_{x}=0$ (A.4)
which, in particular, implies
$\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}_{xx}+\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}_{xx}=-\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}_{x}\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}_{x}-\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}_{x}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}_{x}.$ (A.5)
Thus, the first braces in the right-hand side of equation (A.3) disappear while the factor in front of $u$ becomes just 2uv-2uv. Equation (A.3) now reads
$u_{t}=-u_{xx}-2u^{2}v$ (A.6)
which is nothing but the first equation from (2.26).
The second equation from (2.26) can be demonstrated in the similar way.
Appendix B. Validation of the ansatz.
@@ Acknowledgments.
$\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}}-\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}$ $=$ $\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})|\beta\rangle-\langle{a}|\mathsf{R}^{-1}\mathsf{F}|\beta\rangle$ (B.1)
$=$ $\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})|\beta\rangle-\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}\mathsf{J}_{\xi}\mathsf{F}|\beta\rangle$
@@
$\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}}-\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}=\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{X}\mathsf{F}|\beta\rangle$ (B.2)
@@
$(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{X}\mathsf{F}=\mathbb{E}_{\xi}{\mathsf{F}}-\mathsf{J}_{\xi}\mathsf{F}.$ (B.3)
@@
$\mathsf{X}$ $=$ $1-\mathsf{J}_{\xi}+\mathsf{B}\mathsf{A}-(\mathbb{E}_{\xi}{\mathsf{B}})(\mathbb{E}_{\xi}{\mathsf{A}})\mathsf{J}_{\xi}$ (B.4)
@@
$=$ $\xi\mathsf{R}^{-1}\mathsf{F}^{-1}+\xi\mathsf{R}^{-1}|\beta\rangle\langle b|\mathsf{L}^{-1}\mathsf{A}.$
Substituting this expression into (B.2) and using (3.8) together with the identity $\mathsf{A}\mathsf{F}=\mathsf{G}\mathsf{A}$, one arrives at
$\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}}-\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}$ $=$ $\xi\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{R}^{-1}|\beta\rangle+\xi\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{R}^{-1}|\beta\rangle\langle b|\mathsf{L}^{-1}\mathsf{A}\mathsf{F}|\beta\rangle$ (B.5)
@@
$\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}}-\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}$ $=$ $\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})(\mathbb{E}_{\xi}{\mathsf{B}})|\alpha\rangle-\langle a|\mathsf{R}^{-1}\mathsf{F}\mathsf{B}|\alpha\rangle$ (B.6)
$=$ $\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{Y}\mathsf{G}|\alpha\rangle$
@@
$(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{Y}\mathsf{G}=(\mathbb{E}_{\xi}{\mathsf{F}})(\mathbb{E}_{\xi}{\mathsf{B}})-\mathsf{J}_{\xi}\mathsf{F}\mathsf{B}.$ (B.7)
@@
$\mathsf{Y}$ $=$ $\mathbb{E}_{\xi}{\mathsf{B}}-\mathsf{J}_{\xi}\mathsf{B}+(\mathbb{E}_{\xi}{\mathsf{B}})\mathsf{A}\mathsf{B}-(\mathbb{E}_{\xi}{\mathsf{B}})(\mathbb{E}_{\xi}{\mathsf{A}})\mathsf{J}_{\xi}\mathsf{B}$ (B.8)
@@
$=$ $-\xi\mathsf{R}^{-1}|\beta\rangle\langle b|\mathsf{L}^{-1},$
and substituting it in (B.6) one can obtain
$\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}}-\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}$ $=$ $-\xi\langle b|\mathsf{L}^{-1}\mathsf{G}|\alpha\rangle\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{R}^{-1}|\beta\rangle$ (B.9)